Antireflection glasses. You're applying a 110nm-thick antireflection coating to eyeglasses, hoping to limit reflection in the middle of the visible spectrum (? = 550 nm).

Part A

What should be the coating's refractive index, assuming it's less than that of the lenses?

Part B

If instead you use a coating with n = 1.34 , what's its minimum thickness?
Answer #1

(n-1)*t = lambda/2

or, (n-1)110 = 550/2

so, n= refractive index = 3.5

part B

(1.34-1)*t = 550/2

or, t = 808.82 nm